# disposition

### v7

@@ Abstract
Abstract. This paper develops a formal framework for analyzing information sufficiency in cooperative decision-making. Departing from models of incomplete information that emphasize accuracy of beliefs, we ask what kind and amount of information are required for justifying a choice. Using sequent calculus (Gentzen, 1935a, b), we construct a syntactic model that makes explicit an agent’s reasoning process: from which information, through what inference procedure, and to what conclusion. We define a syntactic criterion, called Ci-acceptability, to determine when a proposed distribution is justifiable given limited structural information. We show that, under a mild assumption that each coalition’s potential is known to at least one member, the set of unanimously accepted payoffs coincides with the core. This reinterprets the core not as a predictive solution, but as the boundary of what can be justified under minimal information. Moreover, our results offer a new perspective on Debreu and Scarf’s (1963) theorem: while the core converges to the competitive equilibrium in replicated markets, this convergence obscures a key asymmetry. Competitive equilibrium requires only local information—preferences and prices—regardless of market size, whereas reaching the core entails increasing informational demands that may ultimately exceed any agent’s cognitive capacity.
@@
Information sufficiency, Cooperative game theory, The core, Debreu-Scarf theorem, Proof theory
@@
Someone saw Nasrudin searching for something on the ground. “What have you lost, Mulla?” he asked. “My key,” said the Mulla. So they both went down on their knees and looked for it. After a time the other man asked: “Where exactly did you drop it?” “In my own house.” “Then why are you looking here?” “There is more light here than inside my own house.”—– “The exploits of the Incomparable Mulla Nasrudin” by Idries Shah, quoted in Doniger (2009)
@@ 1 Introduction
What kind of information—and how much of it—is sufficient for decision-making, particularly for justifying a proposed payoff distribution in a cooperative setting? This fundamental question not only lies at the heart of understanding decision-making under complicated circumstance where relevant information may be missed, it also has profound implications for how we interpret and apply canonical results in economic theory. Consider the classical Debreu-Scarf theorem (Debreu and Scarf, 1963), which demonstrates that the core converges to competitive equilibrium in replicated markets. From the aspect of information sufficiency, however, it obscures a crucial asymmetry: competitive equilibrium requires only local information—individual preferences and market prices—regardless of market size, whereas reaching the core entails increasingly comprehensive information about global cooperation potentials. This asymmetry raises fundamental questions about the theorem’s interpretation and practical applicability, particularly when we consider the cognitive limitations of real decision-makers.
@@
To address this problem, one may consider of game theoretical models with incomplete information. Yet much of the existing literature on decision-making under incomplete information presupposes that players form probabilistic beliefs over a space of possible states. This assumption, deeply rooted in the Harsanyi’s doctrine, enables sophisticated modeling of strategic reasoning and communication. However, it also implicitly imposes a cognitively demanding structure: agents must possess subjective probability measures even when confronting profound uncertainty about the scope and nature of their ignorance.
This paper departs from that probabilistic framework to explore a different fundamental question: under what conditions can an agent justify a decision based solely on the information she actually possesses—without invoking probabilistic beliefs about unknown elements? Like Nasrudin in the epigraph, decision-makers often work with the information that is readily available rather than seeking what might be more relevant, either because they are unaware of what they do not know, or they recognize informational gaps but lack the motivation to fill them. We develop a model to capture this situation, where players possess partial, structural information about the situation, such as the values of specific coalitions, but neither quantify their ignorance through probability assignments nor engage in sophisticated reasoning about uncertainty.
We focus on the core, the canonical solution concept in cooperative games. The core, like other solution concepts, specifies a set of payoff vectors satisfying certain conditions, but the game formulation itself remains silent about what actions players can take and how these actions relate to achieving particular payoff vectors. To bridge this gap between the core and individual decision-making, we transform the cooperative game into a simple decision problem by asking each agent whether she would accept a proposed payoff vector. To give an answer, the agent must make logical inferences from the knowledge she possesses. [^2] The yes-or-no-to-a-proposal model bears similarity to Serrano and Vohra’s (2007). Yet our model does not require agents to engage in strategic sophistication or consider others’ behavior and beliefs. Our central theoretical concern is then to formalize what it means for information to be sufficient for “correct” acceptance with respect to the core. Rather than asking whether players can form optimal beliefs given their information, we ask whether the information they possess can logically support an unanimous acceptance of a proposed payoff vector if and only if the payoff vector is in the core.
This approach stands in marked contrast to existing models of cooperative games with incomplete information that rely on information partitions and belief hierarchies (Wilson, 1978; Serrano and Vohra, 2007; Liu, Mailath, Postlewaite, and Samuelson, 2014; Pomatto, 2022). [^3] Our model is in spirit closer to the model of unawareness (Heifetz, Meier, and Burkhard, 2013) and misspecification (Esponda and Puozo, 2016; Cerreia-Vioglio, Hansen, Maccheroni, Marinacci, 2025). Instead of modeling players’ strategic reasoning, we focus exclusively on one’s exogenous knowledge about the situation, that is, the potential chance of cooperation, without assuming that she has a theory of mind. We adopt this approach to identify the minimal informational skeleton underlying cooperation. By studying what each player can infer solely from her own information about the situation, we isolate the baseline conditions under which a payoff vector is acceptable for her. Such individual assessments form the base for cooperative communication, as cooperation typically begins when someone recognizes a potential Pareto improvement for a group of participants (Olson, 1965).
To define sufficiency of information, we need to determine the unit of information—that is, atomic sentences whose content concerns the decision-making situation. This requires a formal language that can precisely express statements about coalitional values. We modify the framework developed in Bonanno (2008, 2015) and syntactically encode the description of the situation. We model decision-making proof-theoretically by employing a Gentzen (1935a; 1935b)-style sequent calculus modified by Kaneko (2002); Kaneko and Suzuki (2002) to adapt to decision-theoretical analysis. [^4] The use of proof-theoretic methods to study strategic reasoning was initiated by Kaneko and Nagashima (1996, 1997). See Kaneko (2002) for a survey. Here, an agent’s reasoning process is formulated as a proof from ground assumptions (i.e., the information that she possesses) using classical inference rules. The decision criterion is formalized as a statement called Ci-acceptability: a payoff vector is Ci-acceptable for an agent if it cannot be logically refuted given the coalitional potentials known to her. Hence, whether to accept a proposed payoff vector $x$ is determined by whether the agent can prove that $x$ is Ci-acceptable based on her available information.
Our main result demonstrates that, under the minimal requirement that every coalition is known to at least one of its members, the set of payoff vectors that are unanimously Ci-acceptable coincides exactly with the core. Since obtaining and processing information burdens cognitive systems, this finding reinterprets the core not as a behavioral predictive solution but as a cognitive boundary. This reinterpretation illuminates a previously under-appreciated asymmetry in Debreu-Scarf theorem: while the mathematical convergence result remains valid, our analysis reveals that the informational requirements for justifying core outcomes may be prohibitively demanding in practice, suggesting that competitive equilibria—with their modest informational demands—may be more robust in the cognitive sense.
@@
In general, a meaningful economic theory must grapple with cognitive limitations, recognizing that the most elegant theoretical results may be practically unattainable if they demand superhuman cognitive capabilities. By grounding cooperative game theory in logical derivability rather than probabilistic reasoning, this paper opens a new perspective on solution concepts in economics. Our approach builds bridges between individual reasoning, formal logic, and cooperative decision theory, offering a framework that, via information sufficiency, explores the cognitive limitations of real decision-makers while maintaining theoretical rigor. Further, the results provide insight into the minimal informational foundations required for cooperative behavior and suggest new directions for understanding how cooperation emerges under limited information.
@@ 2 TU Games and the Core: An Overview
A central question in cooperative game theory is which payoff vectors should be considered “desirable” or acceptable given the structure of the game. One natural requirement is efficiency, or Pareto optimality, meaning that a payoff vector is a full distribution of the income earned by the comprehensive coalition $N$. Another compelling requirement is coalitional stability, which states that no subset of players should have an incentive to deviate and form a coalition that can secure a better outcome for its members. These two conditions together define the core, a canonical solution concept in cooperative game theory that captures efficient and stable (“fair”) payoff distributions.
@@ 3 Describing decision-making in a Formal Language
Intuitively, a formal language has the following components: [^8] For a thorough introduction formal languages, inference rules, and proofs, see…. Though typically not part of the standard curriculum for economists, some game theorists use models built on formal logical language to study strategic reasoning in games. See, for example, Battigalli and Bonanno (1999), Arieli and Aumann (2015).
@@ 3.3 Formal Language in a Nutshell II: Axioms, Inference Rules, and Proof
$\frac{\Gamma\Rightarrow\Theta,A\ \ \ \ B,\Gamma\Rightarrow\Theta}{A\rightarrow B,\Gamma\Rightarrow\Theta}(\rightarrow\text{-Lef) \ \ \ \ \ \ \ \ \ \ \ }\frac{A,\Gamma\Rightarrow\Theta,B}{\Gamma\Rightarrow\Theta,A\rightarrow B}\text{(}\rightarrow\text{-Right}\text{)}$
@@
The inference rules of sequent calculus are significant because they provide a systematic, symmetric framework where logical operations are broken down into their most basic proof-theoretic components, making the structure of logical reasoning transparent and enabling powerful meta-theoretical results like cut-elimination (see footnote 13), providing a foundation for understanding computation and proof search.
@@ 4 Information Sufficiency for Unanimous Acceptance of Core Payoff Vectors
While condition (D) ensures logical maximal consistency, it is relatively loose from a contextual perspective: a set $\Gamma_{i}$ can satisfy (D) while still encoding information that is contextually unreasonable about coalition formation. For instance, it is possible for $\Gamma_{i}$, which satisfies (D), to include both $y^{S}$ and $\lnot\tilde{y}^{S}$ for some vectors $y,\tilde{y}$ such that $\sum_{j\in S}\tilde{y}_{j}<\sum{j\in S}y_{j}$. In such a case, $\Gamma_{i}$ affirms that $y$ is feasible for $S$ while denying the feasibility of $\tilde{y}$, even though $\tilde{y}$ requires fewer resources than $y$ and should therefore be at least as feasible. This illustrates that $\Gamma_{i}$ may satisfy (D) without conveying any meaningful or reasonable information about coalition values.
@@
For all coalitions outside of $\mathcal{S}$, every possible payoff vector $y^{T}$ is rejected, i.e., $\lnot y^{T}\in\Gamma^{G}(\mathcal{S})$. This captures a strong modeling assumption: when an agent lacks information about a coalition—whether due to unawareness, ignorance, or uncertainty—they systematically deny all payoff possibilities for that coalition. In other words, $\Gamma^{G}(\mathcal{S})$ packages epistemic limitations (such as unawareness or insufficient information) into a syntactic simplification: total denial of feasibility for unknown coalitions. [^15] While this may seem extreme, as shown in the behavior of Nasrudin in the epitaph, this is not abnormal. Neither it is without precedent in the literature. A similar idea appears in decision-theoretic treatments of ambiguity and model misspecification where states outside an agent’s belief model are assigned utility $\infty$ (see Gilboa and Schmeidler, 1989). By relaxing our restriction and introducing ideas from fuzzy logic, one may bridge our model with the literature on incomplete information in cooperative games. Further research is needed in this direction. In what follows, we restrict our attention to information sets of the form $\Gamma^{G}(\mathcal{S})$ for some $\mathcal{S}\subseteq 2^{N}$.
@@
    If $\cup_{i\in N}\mathcal{S}_{i}=2^{N}-\{\emptyset\}$, then for each $G\in\mathbb{G}(N,\cdot)$ and each $x\in\mathbb{N}^{N}$, $x\in C(G)$ if and only if $\vdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow C_{i}(x)$ for all $i\in N$.
@@
    If $\cup_{i\in N}\mathcal{S}_{i}\subsetneq 2^{N}-\{\emptyset\}$, then there is some $G\in\mathbb{G}(N,\cdot)$ and some $x\in\mathbb{N}^{N}$ such that $x\not\in C(G)$ but $\vdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow C_{i}(x)$ for all $i\in N$.

### v8

@@ Abstract
Abstract. This paper develops a formal framework for analyzing information sufficiency in cooperative decision-making. Departing from models of incomplete information that emphasize accuracy of beliefs, we ask what kind and amount of information are required for justifying a choice. Using sequent calculus (Gentzen, 1935a, b), we construct a syntactic model that makes explicit an agent’s reasoning process: from which information, through what inference procedure, and to what conclusion. We define a syntactic criterion, called Ci-acceptability, to determine when a proposed distribution is justifiable based on the information one possesses. We show that if every coalition is known to at least one of its members, then the set of unanimously accepted payoffs coincides with the core; moreover, this informational requirement is minimal when we consider all games sharing the same set of players. This reinterprets the core not as a predictive solution, but as the boundary of what can be justified under minimal information. Moreover, our results offer a new perspective on Debreu and Scarf’s (1963) theorem: while the core converges to the competitive equilibrium in replicated markets, this convergence obscures a key asymmetry: competitive equilibrium requires only local information—preferences and prices—regardless of market size, whereas reaching the core entails increasing informational demands that may ultimately exceed any agent’s cognitive capacity.
@@
Information sufficiency, Cooperative game theory, The core, Debreu-Scarf theorem, Formal language, Proof theory
@@
Someone saw Nasrudin searching for something on the ground. “What have you lost, Mulla?” he asked. “My key,” said the Mulla. So they both went down on their knees and looked for it. After a time the other man asked: “Where exactly did you drop it?” “In my own house.” “Then why are you looking here?” “There is more light here than inside my own house.”
—– Idries Shah, The Exploits of the Incomparable Mulla Nasrudin
“If three people drink three bottles of water in three days, how many bottles of water will nine people drink in nine days?” “Nine.”
— A widely shared riddle on social media and its most common answer
Your faith was strong, but you needed a proof.
— Leonard Cohen, “Halleluja”
@@ 1 Introduction
A central topic of classical economic theory is the relationship between information and behavioral outcomes. The dominant approach in the literature has been to examine how available information —albeit incomplete — can be used effectively, and what constitutes a rational decision given such information. However, this perspective leaves a complementary question under-explored: what kinds of information are sufficient for efficient decision-making, and what decisions will a particular information set necessarily produce? This inverse perspective is central to understanding behavior in complex circumstances, where relevant information may be limited by physical or cognitive constraints. Exploring these questions can help illuminate canonical results in economic theory and provide insights into the functioning of economic, political, and social systems.
Consider, for instance, the classical Debreu and Scarf’s (1963) theorem, which demonstrates that under replication, the core shrinks to the competitive equilibrium set. Yet this asymptotic identity obscures a crucial asymmetry: competitive equilibrium requires only local information — individual preferences and market prices — regardless of market size, whereas reaching the core entails increasingly comprehensive information about global cooperation potentials, and the informational requirements therefore grow with the number of agents.
Once this asymmetry of information sufficiency is taken seriously, the usual interpretation of the convergence theorem loses its force. Initiated by Jevons (1871) and Edgeworth (1881), the result is often taken to justify competitive equilibrium as an appropriate model for large economy since it is the limit of decentralized bargaining. Yet if agents have bounded cognitive or informational capacities, they may be unable to identify or verify core allocations. In this sense, the classical convergence result may fail behaviorally even if it holds mathematically, weakening one of the canonical micro-foundations for competitive equilibrium.
A more constructive question raised by this asymmetry is: what kinds of information are sufficient for reaching efficient outcomes? The core is attractive precisely because its allocations are both stable and efficient — no coalition can improve upon them through re-contracting. When relevant information is fully available, stability and efficiency coincide in this sense. But when some information is inaccessible, an allocation may remain stable merely because no coalition can discover a profitable deviation, even though it is inefficient from an outside perspective. This separation suggests that stability achieved under informational constraints may be systematically weaker than efficiency. It also raises a further question: how does the sufficient information required to restore this equivalence grow as the economy becomes large?
The literature on the core under incomplete information, beginning with Wilson (1978), offers important guidance for this perspective. [^2] See Vohra (1999); Dutta and Vohra (2005); Serrano and Vohra (2007); de Clippel (2007); Myerson (2007); Serrano and Vohra (2007), among others, for detailed discussion. Also, for studies on incomplete information and stable matching, which is a closely related topic, see, among others, Roth (1989); Chade (2006); Ehlers and Massó (2007); Hoppe, Moldovanu, and Sela (2009); Liu, Mailath, Postlewaite, and Samuelson (2014); Pomatto (2022); Chade, Lewis, and Smith (2014). In these models, a cooperative game is reformulated as an explicit decision problem for each agent: a doxastic/epistemic structure representing her beliefs, a set of feasible actions, and a mapping from action profiles to outcomes. This reformulation transforms the cooperative setting into a decentralized problem of individual strategic reasoning. A central insight of this literature is that stability under incomplete information generally diverges from efficiency under complete information. That is, solution concepts based on agents’ limited knowledge may yield outcomes that are stable under incomplete information framework yet fail to coincide with the efficient core.
@@
Yet this literature does not address the question raised here. Models of incomplete information focus on how accurate agents’ information is, not what kinds of information — or how much of it — are sufficient for making cooperative decisions. In the standard Harsanyi framework, players are assumed to hold probabilistic beliefs over the entire space of possible states. This assumption allows for rich representations of strategic reasoning and communication, but it also imposes a cognitively demanding structure: agents must entertain well-defined subjective probabilities even when confronting profound ignorance about the scope and nature of their ignorance.
Our question is different: if an outcome is to be justified solely on the basis of the information agents actually possess, then what conditions must that information satisfy for only efficient outcomes to be justifiable? Agents are not required to form probabilistic beliefs about unknown or unrecognized contingencies. What matters instead is that they understand and focus on the information available to them that they can rely upon, and try to imply useful conclusion from it, for example, they may set a well-defined criterion for what it means for an outcome to be “acceptable,” and reason — using only their own information — if a proposed outcome meets this criterion.
We follow Serrano and Vohra (2007) in recasting a cooperative game as a collection of decentralized decision problems, where each agent must decide whether to accept a proposed payoff vector. Our approach departs from theirs, however, by focusing on how agents navigate from information to decision. Therefore, instead of assuming sophisticated strategic reasoning, we assume that each agent performs only logical inference based on the information she actually possesses.
This perspective requires an explicit representation of information — one that is both versatile and quantifiable. It must be versatile in the sense of being syntactically flexible and grammatically independent of meaning, so that it can encode a wide range of information relevant for decision-making. It must also be quantifiable, allowing us to measure the size of an agent’s information and categorize its content. Moreover, we need a tractable and constructive account of an agent’s reasoning procedure — one that makes the assumed cognitive abilities explicit rather than relying on implicit semantic knowledge, such as classical set-theoretic computation.
To provide a framework that meets these requirements, we adopt a purely syntactic approach. Our modeling proceeds in three steps. First, we give an explicit definition of information. Drawing inspiration from Bonanno’s (2008; 2015) formally linguistic encoding of non-cooperative games, we construct a formal language that precisely expresses statements about coalitional feasibility and domination. These syntactic representations encode the situation’s description and form the information foundation for decision-making in our context.
Second, within this language, we introduce the decision criterion, a statement termed Ci-acceptability. Informally, a payoff vector $x$ is Ci-acceptable for agent $i$ if there does not exist any coalition $S$ including $i$ and any payoff vector $y$ such that (1) $y$ is feasible for $S$, (2) $y$ weakly dominates $x$, and (3) the $i$-th coordinate of $y$ is strictly larger than that of $x$. One can see that Ci-acceptability individualizes (or decentralizes) the criteria embodied in the core: it faithfully rephrases the definition of the core, and condition (3) emphasizes the incentive for agent $i$ to reject the proposed vector $x$.
With this well-defined criterion, the mechanism of decision-making becomes clear-cut: an agent accepts a payoff vector if she can verify that the vector is Ci-acceptable and rejects it if it is not. We then formalize the agent’s verification process as a proof. Much like branches of rivers converge and ultimately flow into the ocean, a proof is a tree-structured sequence of statements that begins with axioms and proceeds via inference rules. To articulate these proofs, we adopt a version of the classic Gentzen-style sequent calculus (Gentzen, 1935a, b) modified by Kaneko (2002). This system offers an intuitive and tractable language for describing the flow of proof. Its inference rules encode the most basic logical reasoning capacities shared by all models of rational decision-making, without appealing to set-theoretic or probabilistic assumptions. Equipped with this proof system, we can trace precisely which information is needed to establish any given statement.
Our main question concerns the conditions under which efficiency coincides with stability: under what informational conditions is a payoff vector in the core if and only if no agent rejects it based on the information she possesses? Under a mild technical simplification that ensures logical decidability, this equivalence can be restated as the condition under which a payoff vector is in the core if and only if it is unanimously accepted (Proposition 4.3).
To approach this question gradually, we categorize information according to coalitions and study how the amount and distribution of information among agents affect the coincidence of efficiency and stability. Proposition 4.4 provides a starting point: any payoff vector in the core is always unanimously accepted, regardless of the information structure. In the extreme case, even if agents have no information at all, such a payoff vector will not be rejected by anyone. Thus, the essential challenge is to determine when a non-core payoff vector will be rejected by at least one agent.
It is natural to expect that agents require more information to reject a payoff vector than to accept it. Our first auxiliary result formalizes part of this requirement. Observation 1 shows that information about coalitions that do not include an agent is irrelevant for her decision-making. To reject a payoff vector, an agent needs only information about coalitions in which she herself participates.
Our second auxiliary result identifies an extreme informational condition under which efficiency and stability do coincide. Observation 2 states that if every agent possesses information about all coalitions that include her, then efficiency is equivalent to unanimous acceptability. This provides an upper bound on the amount of information sufficient for the equivalence. The remaining question is whether such a large information requirement is truly necessary: does every agent really need to know that much?
Our main result, Theorem 4.5, shows that when every coalition is known to at least one of its members, the set of payoff vectors that are unanimously Ci-acceptable coincides exactly with the core. Moreover, the theorem establishes that this informational requirement is minimal when we consider the class of all games with the same set of agents: if even a single coalition is not known to any of its members, then there exists a game in which some non-core payoff vector fails to be rejected by any agent.
We then apply these ideas to the Debreu-Scarf theorem. Focusing on the simple pure-exchange replica economy studied in Shapley (1975), we adapt our formal language to this environment, relying in particular on the categorization of coalition-specific information. Our analysis shows that as the number of replicas increases, the average informational burden required to reject non-core allocations — and thereby to approach the unique competitive equilibrium — grows rapidly. In the limit, this burden exceeds any reasonable bound on an individual agent’s cognitive capacity.
This perspective formally reveals a previously under-appreciated asymmetry in the Debreu-Scarf theorem: although the mathematical convergence result remains intact, the informational requirements needed for agents to justify core outcomes may be prohibitively demanding in practice. Our results therefore suggest that the core — or any solution concept — should be examined not merely as a normative benchmark but also in light of the cognitive constraints agents face, before being treated as an appropriate behaviorally predictive solution.
@@
In general, a meaningful economic theory must grapple with cognitive limitations, recognizing that the most elegant theoretical results may be practically unattainable if they demand superhuman cognitive capabilities. By grounding cooperative game theory in logical derivability, our paper opens a new perspective on solution concepts in economics. Our approach builds bridges between individual reasoning, formal logic, and cooperative decision theory, offering a framework that, via information sufficiency, explores the cognitive limitations of real decision-makers while maintaining theoretical rigor. Further, the results provide insight into the minimal informational foundations required for cooperative behavior and suggest new directions for understanding how cooperation emerges under limited information.
Finally, one may ask why, given that other frameworks—such as models of unawareness (Heifetz, Meier, and Burkhard, 2013)—address questions similar in spirit to ours, we choose to adopt a formal language and a proof-theoretic approach. Our first reason is conceptual. Human beings store information and reason through language, so it is natural to model both information and its processing using a formal language grounded in symbolic logic. This perspective is not new. The usefulness of formal languages and logical tools is well established in economic theory. [^3] See Gonczarowski, Kominers, and Shorrer (2025) for a recent example, which employs a formal language to formulate the model and uses the compactness theorem in logics to scale finite results to infinite ones. In fields directly connected to doxastic and epistemic considerations — most notably epistemic game theory — formal languages and logical methods have long been central. [^4] See Fagin, Halpern, Moses, and Verdi (1995); Battigalli and Bonanno (1999); Bonanno (2015) for surveys.
By contrast, proof theory has been used far less frequently. [^5] The use of proof-theoretic methods to analyze strategic reasoning originates with Kaneko and Nagashima (1996, 1997). See Kaneko (2002) for a survey. The literature typically relies on semantic, set-theoretic models of reasoning. We turn instead to a proof-theoretic, constructive approach because it makes the reasoning process explicit. Proofs literally describe the flow of thought, exposing the internal structure of reasoning in a way that semantic models often obscure. Moreover, a purely syntactic account avoids inadvertently embedding assumptions about agents’ cognitive abilities — assumptions that semantic models may impose without acknowledging them. This feature makes proof theory well suited to represent various forms of reasoning, including reasoning guided by shallow linguistic heuristics rather than mathematical analysis, as exemplified by the answer to the epigraph’s puzzle.
Ultimately, while many tools can be used to organize and visualize reasoning, providing an explicit flow of sentences to support a conjecture aligns naturally with our objectives. As Leonard Cohen’s lyrics in the epigraph remind us, at the end of the day, we need a proof.
@@ 2 TU Games and the Core: An Overview
A central question in cooperative game theory is which payoff vectors should be considered acceptable or “desirable” given the structure of the game. One natural requirement is efficiency, or Pareto optimality, meaning that a payoff vector is a full distribution of the income earned by the comprehensive coalition $N$. Another compelling requirement is coalitional stability, which states that no subset of players should have an incentive to deviate and form a coalition that can secure a better outcome for its members. These two conditions together define the core, a canonical solution concept in cooperative game theory that captures efficient and stable (“fair”) payoff distributions.
@@ 3 Describing decision-making in a Formal Language
Intuitively, a formal language has the following components: [^9] For a thorough introduction formal languages, inference rules, and proofs, seeextbooks such as Ebbinghause, Flum, and Thomas (2021) and Mancosu, Galvan, and Zach (2021).
@@ 3.3 Formal Language in a Nutshell II: Axioms, Inference Rules, and Proof
$\frac{\Gamma\Rightarrow\Theta,A\ \ \ \ B,\Gamma\Rightarrow\Theta}{A\rightarrow B,\Gamma\Rightarrow\Theta}(\rightarrow\text{-Left) \ \ \ \ \ \ \ \ \ \ \ }\frac{A,\Gamma\Rightarrow\Theta,B}{\Gamma\Rightarrow\Theta,A\rightarrow B}\text{(}\rightarrow\text{-Right}\text{)}$
@@
The inference rules of sequent calculus are significant because they provide a systematic, symmetric framework where logical operations are broken down into their most basic proof-theoretic components, making the structure of logical reasoning transparent and enabling powerful meta-theoretical results like cut-elimination (see footnote 14), providing a foundation for understanding computation and proof search.
@@ 4 Information Sufficiency for Unanimous Acceptance of Core Payoff Vectors
While condition (D) ensures logical maximal consistency, it is relatively loose from a contextual perspective: a set $\Gamma_{i}$ can satisfy (D) while still encoding information that is contextually unreasonable about coalition formation. For instance, it is possible for $\Gamma_{i}$, which satisfies (D), to include both $y^{S}$ and $\lnot\tilde{y}^{S}$ for some vectors $y,\tilde{y}$ such that $\sum_{j\in S}\tilde{y}_{j}<\sum_{j\in S}y_{j}$. In such a case, $\Gamma_{i}$ affirms that $y$ is feasible for $S$ while denying the feasibility of $\tilde{y}$, even though $\tilde{y}$ requires fewer resources than $y$ and should therefore be at least as feasible. This illustrates that $\Gamma_{i}$ may satisfy (D) without conveying any meaningful or reasonable information about coalition values.
@@
For all coalitions outside of $\mathcal{S}$, every possible payoff vector $y^{T}$ is rejected, i.e., $\lnot y^{T}\in\Gamma^{G}(\mathcal{S})$. This captures a strong modeling assumption: when an agent lacks information about a coalition—whether due to unawareness, ignorance, or uncertainty—they systematically deny all payoff possibilities for that coalition. In other words, $\Gamma^{G}(\mathcal{S})$ packages epistemic limitations (such as unawareness or insufficient information) into a syntactic simplification: total denial of feasibility for unknown coalitions. [^16] While this may appear extreme, as illustrated by Nasrudin’s reasoning in the epigraph, our use of negation is merely a technical simplification. Instead of inserting the negation of a statement into the antecedent, one could allow a set of sentences simply to be absent. Under this alternative, certain cases lead to undecidability of the criterion, which may be reflected behaviorally, for example, by a prolonged silence. If we admit only those outcomes that are unanimously and explicitly accepted, then such a silence is behaviorally indistinguishable from a rejection. In what follows, we restrict our attention to information sets of the form $\Gamma^{G}(\mathcal{S})$ for some $\mathcal{S}\subseteq 2^{N}$.
@@
    If $\cup_{i\in N}\mathcal{S}_{i}=2^{N}-\{\emptyset\}$, then for each $G\in\mathbb{G}(N,\cdot)$ and each $x\in\mathsf{PV}(G)$ feasible for $N$, $x\in C(G)$ if and only if $\vdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow C_{i}(x)$ for all $i\in N$.
@@
    If $\cup_{i\in N}\mathcal{S}_{i}\subsetneq 2^{N}-\{\emptyset\}$, then there is some $G\in\mathbb{G}(N,\cdot)$ and some $x\in\mathsf{PV}(G)$ feasible for $N$ such that $x\not\in C(G)$ but $\vdash\Gamma^{G}(\mathcal{S}_{i})\Rightarrow C_{i}(x)$ for all $i\in N$.